# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

### v3

@@ 3.2. Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG system to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM, it calculates relevance score $R(q,d)$ for each document $d$ in $D$ and returns the set of documents with the $k$ highest relevance score as the retrieval result (the LLM only takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)=\{d_{1},d_{2},...,d_{k}\}$. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$.
@@ 3.2.1. Objective of the Adversary
The attacker aims to find the adversarial trigger $p_{\text{adv}}$ for each $N$ selected target document $d_{t}$, and add it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $R(q,d_{t}\oplus p_{\text{adv}})$ assigned by the retrieval model RM, $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)=\{d\mid d_{t}\oplus p_{\text{adv}}\}$ , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\{d\mid d_{t}\oplus p_{\text{adv}}\}))=S_{t}$, $S_{t}\neq S_{o}$.
@@ 3.2.2. Capabilities of the Adversary
In the black-box scenario, we assume the attacker is only authorized to query the RAG, obtain RAG-generated results and modify at most $N$ documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the attacker has no knowledge of the model architecture, model parameters, or any other information related to the models within the black-box RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited.
@@
Figure 3. The framework for obtaining ranking imitation data of retrieval model in black-box RAG system.
3.3. Black-box Retriever Imitation
The primary issue in implementing manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to imitate the black-box retrieval model $RM$ in RAG. The basic idea is to train a surrogate model $M_{\text{i}}$ with the ranking results $RM_{\text{k}}(q)$ from the retrieval model $RM$, thus turning the black-box retrieval model into a white-box model. However, since the retriever and the LLM in RAG are serially connected, it is not possible to directly obtain the ranking results of the retriever. Therefore, we attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents irrelevant to the query can be randomly sampled as negative examples $d_{-}$ Therefore, we design specific instructions to induce the black-box RAG to replicate the retrieval results of the retriever RM. These retrieval results only need to reflect the relevance to the query. Then, based on the generated results of the LLM and the corpus, we sample positive and negative data to train the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG is illustrated in Figure 3. The specific instruction used is as follows:
@@ 3.3. Black-box Retriever Imitation
—Here is the user question— < < <{query}> > > —Here is the user command— Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context.
After that, we use a pairwise approach to sample data and train the surrogate model. The black-box RAG responds to the instruction with context information so the responses generated by the black-box RAG reflect the retrieval results. Relevant documents are sampled from the responses as positive examples $d_{+}$, and random irrelevant documents are sampled as negative examples $d_{-}$. For real-world RAG system attack, negative samples can be constructed using web information randomly sampled from search engines like Google to simulate real-world adversarial noise in RAG corpora.
@@
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Let the relevance score calculated by $M_{i}$ be $R_{i}$, we train the surrogate model $M_{i}$ by minimizing the following contrastive loss, enabling the surrogate model $M_{i}$ to approximate the internal mechanisms of the black-box RAG retrieval model.:
(1) $L=-\frac{1}{|Q|}\sum_{q\in Q}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$
@@ 3.4. Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) (Liu et al., 2023b) strategy for adversarial retrieval attacks, which is commonly used as the baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its generated optimization function being (Liu et al., 2023b):
(2) $\max\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot\log P_{g}(T_{\text{pat}};w)\\ +\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}$
In the above formula, $P_{g}$ is the semantic constraint function, and $f_{\text{nsp}}$ is the next sentence prediction consistency score function between $T_{\text{pat}}$ and $d_{t}$. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $w$ represents the internal weights of the corresponding model.
Subsequently, we insert $p_{\text{adv}}$ into the target document to execute the poisoning attack on the RAG corpus. Our objective of RAG opinion manipulation attack is as follows:
@@
(4) $\begin{split}s.t.RM(q)&=E(q)\cdot E(d),d\in D(d;d_{t}\oplus p_{\text{adv}})\\ RM_{k}(q)&=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q)\}\end{split}$
I(·) is the indicator function whose return value is 1 if the conditions within the parentheses are satisfied; otherwise, the return value is 0. S(·) is the opinion classification function. The retrieval model computes semantic relevance score by dot product, and $E$ denotes the word embedding vector. $RM_{k}(q)$ denotes the set composed of the top $k$ documents with the highest relevance scores collected from the ranking of black-box retrieval model $RM$.
@@ 4.1. Dataset
The database of the black-box RAG system in our study consists of MS MARCO Passages Ranking dataset (Nguyen et al., 2016) and PROCON.ORG data [^1] https://testing-www.procon.org/, containing about 500 queries and 500,000 documents. We utilize queries in MSMARCO to obtain the source data for guiding the black-box RAG to generate relevant passages (Zhong et al., 2023) to sample data pairs to train the surrogate model. We use topic queries in PROCON.ORG to get controversial topics arguments for opinion manipulation.
@@ 4.2. Experiment Details
We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain [^2] https://python.langchain.com/v0.2/docs/tutorials/qa_chat_history/. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama3), Vicuna-13b-v1.5 (Vicuna) and Mixtral-8x7B-Instruct-v0.1 (Mixtral). They perform well across various tasks among all open-source LLMs. The prompt in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain.
We select three representative dense retrieval models, Contriever, Co-Condenser and ANCE, as the target retrieval model (Gao and Callan, 2021; Izacard et al., 2021; Xiong et al., 2020). By convention, we use dot product between the embedding vectors of queries and candidate documents as their similarity score $R$ in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset.
For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial triggers $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3. In this experiment, the number of poisoned documents for FlippedRAG ranges from 5 to 10.
@@
We adopt Qwen2.5-Instruct-72B as the opinion recognition model. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text. During black-box imitation training, we set the batch size as 256, the number of epoch as 4 and the learning rate as 4e-5. In the process of implementing PAT to generate adversarial triggers, we set the number of beams as 30, the temperature value as 0.4, the learning rate as 0.1 and the sequence length as 15. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.

### v4

@@ 3.2. Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM in RAG, it calculates relevance score $RM(q,d)$ for each document $d$ in $D$ and return the set of documents with the $k$ highest relevance score as the retrieval result to the LLM (the LLM takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)$. $RM(q,d)$ and $RM_{k}(q)$ are formally defined as:
(1) $\begin{split}RM(q,d)&=E(q)\cdot E(d)\\ RM_{k}(q)&=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q,d)\}\end{split}$
$E$ denotes the word embedding vector. The semantic relevance scores are computed by dot product. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the original opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$. S(·) is the opinion classification function.
To alter $S_{o}$ of the black-box RAG, the adversary achieves adversarial manipulation by setting traps for targets. The adversary first identifies controversial topics for manipulation and configures tailored triggers into a small number of documents. When users query these topics, the system retrieves the adversarial documents, causing the RAG model to generate biased outputs, thereby achieving the adversary’s goal of luring users into the trap. This trap-setting methodology can be conceptually analogous to black-hat SEO techniques.
@@ 3.2.1. Objective of the Adversary
The adversary aims to find the adversarial trigger $p_{\text{adv}}$ for each selected $N$ target-doc $d_{t}$, and add it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $RM(q,d_{t}\oplus p_{\text{adv}})$ , $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)^{\prime}$ ($d_{t}\oplus p_{\text{adv}}\in\text{RM}_{k}(q)^{\prime}$) , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\text{RM}_{k}(q)^{\prime}))\rightarrow S_{t}$, $S_{t}\neq S_{o}$.
@@ 3.2.2. Capabilities of the Adversary
In the black-box scenario, we assume the adversary is only authorized to query the RAG-like system to obtain results and modify at most $N$ documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the adversary has no knowledge of the architecture, parameters, or any other information related to the models in RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited. Targeted traps necessitate adversary selection of controversial topics for manipulation, so adversaries possess awareness of user query intents input to the target retriever in RAG.
Following (zou2024poisonedrag; chaudhari2024phantom; cheng2024trojanrag; shafran2024machine), FlippedRAG presents realistic threats to RAG-like systems using data from public platforms that allow user edits. Adversaries can compromise RAG output objectivity by injecting malicious modifications into knowledge sources including Wikipedia, user-generated content (UGC) platforms, and other publicly accessible web resources. As established in studies (carlini2024poisoningwebscaletrainingdatasets; luo2025unsafellmbasedsearchquantitative; lin2024mawseo), the retrieval corpus consists of millions of texts from the Internet, enabling attackers to insert maliciously crafted web content. This demonstrates FlippedRAG’s viable threat potential against real-world black-box RAG systems.
@@
Figure 3. The framework for obtaining ranking imitation data of retrieval model and constructing contrastive pairs in black-box RAG.
3.3. Black-box Retriever Imitation in RAG
The primary issue in implementing retrieval manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to imitate the black-box retrieval model $RM$ in RAG.
Inspired by (liu_order-disorder_2023; wu2023prada), we intend to train a surrogate model $M_{\text{i}}$ with contrastive leaning, thus turning the black-box retrieval model into a white-box model. Contrastive learning represents a widely adopted training paradigm, simple but effective, as shown in (chen2020simpleframeworkcontrastivelearning; he2020momentumcontrastunsupervisedvisual). Its efficacy critically depends on the quality of contrastive sample pairs. Previous works (liu_order-disorder_2023; wu2023prada) on attacking the black-box retrieval model assuming that the adversary has access to the ranking lists of the target model. Leveraging this capability, they perform targeted sampling to obtain high-quality contrastive data pairs, which help them train surrogate models that accurately imitate the ranking preferences of the target black-box model.
However, since the retriever and the LLM in RAG are serially connected, it is unfeasible to directly obtain the pseudo-relevant training data from ranking results of the retriever in RAG, which hinders the construction of effective contrastive samples.
Building upon observations from our pilot study on ChatGPT, we propose a novel methodology for constructing high-quality contrastive pairs tailored for contrastive learning in RAG, leveraging the instruction following capability of LLMs. We attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed most relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents less relevant to the query can be randomly sampled as negative examples $d_{-}$ . This contrastive sample construction methodology effectively captures the ranking preferences of the target black-box retriever in RAG, enhancing the imitation effectiveness of contrastively trained surrogate models. Subsequent empirical validation in black-box imitation experiments confirms the efficacy of our contrastive learning design.
Specifically, we design specific instructions to induce the black-box RAG to replicate the retrieval results of the retriever RM. Then, based on the generated results of the LLM and the corpus accessible under the defined threat model, we use a pairwise approach to sample positive and negative data as training data of the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG and construct contrastive pairs is illustrated in Figure 3. The specific instruction used is as follows:
@@ 3.3. Black-box Retriever Imitation in RAG
—Here is the user question— < < <{query}> > > —Here is the user command— Please COPY all the given context altogether in $[$ ]] including all marks and symbols. Do not omit any sentence of the context.
The black-box RAG responds the instruction with context information, so its generated outputs inherently reflect the underlying retrieval results. Positive samples $d_{+}$ can be sampled from the responses as relevant documents, while negative samples $d_{-}$ can be derived from two distinct sources:
1) Documents from the lower-ranked positions in the RAG context induced by the above instruction can serve as hard negative samples.
(2) Random sampling. According to the defined threat model, a subset of negative samples can be obtained through random sampling from the adversarially targeted corpus, which is publicly accessible and editable, such as search engines, UGC platforms, and web pages.
@@
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset $\mathcal{D}$. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Then we train the surrogate model $M_{i}$ by minimizing the following contrastive loss, enabling the surrogate model $M_{i}$ to approximate the internal mechanisms of the black-box RAG retrieval model:
(2) $L=-\frac{1}{|\mathcal{D}|}\sum_{\small(d_{+},d_{-})\in\mathcal{D}}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$
We employ the Adam optimizer (adam2014method) to minimize the contrastive loss. This objective loss enables the surrogate model to capture the ranking preferences of the target black-box retriever from contrastive sample pairs. Especially, the hard negatives introduced during sample pair construction provide semantically proximate comparisons, significantly enhancing the surrogate model’s approximation fidelity to the target model. The complete black-box imitation procedure is detailed in Phase 1 of Algorithm A.1.
@@ 3.4. Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into the manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) (liu_order-disorder_2023) strategy for adversarial retrieval attacks, which is commonly used as a baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its loss function being:
(3) $\max_{w}\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot$ $\log P_{g}(T_{\text{pat}};w)+\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}$
where $P_{g}$ is the semantic constraint with a language model $g$. $f_{\text{nsp}}$ is the next sentence prediction consistency score function between $T_{\text{pat}}$ and $d_{t}$ with a next sentence prediction model. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters with value range bounded within the interval [0, 1] inclusively. $w$ represents the internal weights of the corresponding model. We employ stochastic gradient descent to jointly optimize the dual objectives of relevance enhancement and stealth constraint, thereby generating adversarial triggers that both improve relevance rankings and remain imperceptible to humans. The complete algorithm of PAT is detailed in (liu_order-disorder_2023).
Subsequently, we insert $p_{\text{adv}}$ into the target document with $S_{t}$ to execute the poisoning attack on the RAG corpus. $p_{\text{adv}}$ can elevate the ranking position of target documents in RAG, allowing their inclusion in the augmented context provided to the LLM. Thus, we can steer the LLM in the RAG system to generate outputs systematically aligned with the target opinion $S_{t}$. We formally define our attack objective as follows:
@@
I(·) is the indicator function whose return value is 1 if the conditions within the parentheses are satisfied; otherwise, the return value is 0.
To achieve this adversarial objective (4), we should:
1) Optimize contrastive loss (2) during the black-box imitation phase to derive a surrogate model approximating the ranking preferences of the target retrieval model;
2) Optimize the loss function (3) by searching adversarial triggers that simultaneously enhance relevance rankings and maintain stealth properties.
These mechanisms collectively maximize the retrieval rank elevation of documents expressing the target opinion, thereby increasing their proportional presence in the LLM’s context. This cascading effect ultimately amplifies the likelihood that RAG-generated responses align with the target opinion. The opinion manipulation procedure is detailed in Phase 2 of Algorithm A.1.
@@ 4.1. Dataset
The database of the black-box RAG in our study consists of a subset of MS MARCO Passages Ranking dataset (nguyen2016ms) and PROCON.ORG data [^1] https://testing-www.procon.org/, containing about 500 queries and 500,000 documents. We utilize queries in MSMARCO to obtain the source data for guiding the black-box RAG to generate relevant passages (zhong2023poisoning) to sample data pairs to train the surrogate model. We use topic queries in PROCON.ORG to get controversial topics arguments for opinion manipulation.
@@ 4.2. Experiment Details
We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain [^2] https://python.langchain.com/v0.2/docs/tutorials/qa_chat_history/. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama-3), Vicuna-13b-v1.5 (Vicuna) and Mixtral-8x7B-Instruct-v0.1 (Mixtral). They perform well across various tasks among all open-source LLMs. The prompt in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain.
We select four representative dense retrieval models, Contriever, Co-Condenser, ANCE and Qwen3-Embedding-4b [^3] https://huggingface.co/Qwen/Qwen3-Embedding-4B, as the target retrieval model (gao2021unsupervised; izacard2021unsupervised; xiong2020approximate). Qwen3-Embedding-4B represents the latest proprietary model in the Qwen series, ranking among the top three high-performance embedding models on the MTEB leaderboard (Embedding Leaderboard) [^4] https://huggingface.co/spaces/mteb/leaderboard. By convention, we use dot product between the embedding vectors of queries and candidate documents as their similarity score in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset. We selected this model as the surrogate following (liu_order-disorder_2023), wherein NBoost/pt-bert-base-uncased-msmarco demonstrated robust imitation capabilities. This choice is further justified by its domain-specific pre-training for retrieval ranking tasks and compact architecture facilitating efficient fine-tuning.
For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial trigger $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ . In this paper, $k$ is set to 3. In this experiment, the number of poisoned documents for FlippedRAG ranges from 3 to 10. The normal performance of RAG is in Table 14 in the Appendix. Benchmarked against vanilla RAG performance reported in FlashRAG(jin2025flashrag) and TrustRAG(zhou2025trustragenhancingrobustnesstrustworthiness), our implemented RAG achieves comparable efficacy.
@@
We adopt Qwen2.5-Instruct-72B as the opinion recognition model. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text. During black-box imitation training, we set the batch size as 256, the number of epoch as 4 and the learning rate as 4e-5. In the process of implementing PAT to generate adversarial triggers, we set the number of beams as 30, the temperature value as 0.4, the learning rate as 0.1 and the sequence length as 15. All our experiments are conducted on a NVIDIA DGX H100 GPU.